I have a question...its a word problem leading to a quadratic equation
The dimensions of a right angle triangle are given in terms of x.2x is the hypotenuse,(x+3) is the opposite and (x+1) is the adjacent.find x

Answer by mananth(16946)

By Pythagoras Theorem
(x+3)^2+(x+1)^2=(2x)^2

x^2+6x+9+x^2+2x+1=4x^2
2x^2+8x+10=4x^2
2x^2-8x-10=0
/2
x^2-4x-5=0
x^2-5x+x-5=0
x(x-5)+1(x-5)=0
(x-5)(x+1)=0
x= 5 OR x=-1
ignore negative
x= 5